How do I solve this problem?3. Skips provides repair services (R) using two inputs: labor (L), and tools (T). Skips production function is R-L4T1/4. Sk
Answer #1

Total Revenue=TR=P*R=100L1/4T1/4

Total Cost=TC=wL+rT=1*L+5T=L+5T

Profit=TR-TC=100L1/4T1/4-L-5T

a)

We have derived that

Profit=100L1/4T1/4-L-5T

b)

R=L1/4T1/4

Put T=16

R=L1/4161/4=2L1/4

Marginal Product of labor=MPL=dR/dL=2(1/4)L-3/4=0.50L-3/4

Marginal Revenue Product of labor=MRPL=P*MPL=100*0.50L-3/4

Set MRPL=w for profit maximization

100*0.50L-3/4=1

50L-3/4=1

c)

We have derived that

50L-3/4=1

L3/4=50

L=184.20

d)

Profit=TR-TC=100L1/4T1/4-L-5T

d(Profit)/dL=100*(1/4)L-3/4T1/4-1

d(Profit)/dT=100*(1/4)L1/4T-3/4-5

Set d(Profit)/dL=0 and d(Profit)/dT=0

100*(1/4)L-3/4T1/4-1=0

25L-3/4T1/4=1--------------------(1)

100*(1/4)L1/4T-3/4-5=0

25L1/4T-3/4=5 ---------------------(2)

e)

Divide equation 1 by equation 2, we get

(25L-3/4T1/4)/(25L1/4T-3/4)=1/5

(T/L)=(1/5)

L=5T

From equation 2, we get

25L1/4T-3/4=5

i.e.

L1/4T-3/4=5/25=0.20

Put L=5T

(5T)1/4T-3/4=0.20

51/4T1/4T-3/4=0.20

0.20T1/2=51/4

T1/2=5*51/4

T=25*51/2=55.9017

L=5T=5*55.9017=279.5085